### v4

# Sequential Solution Concepts in Cooperative Games with Generalized Characteristic Functions

Abstract
Motivated by the fact that the worth of a coalition may depend on the order in which agents arrive, Nowak and Radzik (1994) (NR) introduced cooperative games with generalized characteristic functions. We study such temporal cooperative games (TCGs), where the worth function $v$ is defined on sequences of agents $\pi$ rather than sets $S$. This order sensitivity necessitates a re-examination of axioms for reward sharing. NR and subsequent work proposed several axioms; the resulting solution concepts are still inherently order-oblivious and closely tied to the Shapley value.
In contrast, we focus on sequential solution concepts that explicitly depend on the realized order $\pi$. We study reward-sharing mechanisms satisfying incentive for optimal arrival (I4OA), which promotes orders maximizing total worth; online individual rationality (OIR), which ensures agents are not harmed by later arrivals; and sequential efficiency (SE), which requires that the worth of any sequence is fully distributed among its agents. These axioms are intrinsic to TCGs, and we characterize a class of reward-sharing mechanisms uniquely determined by them. The classical Shapley value does not directly extend to this setting. We therefore construct natural Shapley analogs in two worlds: a sequential world, where rewards are defined for each sequence–agent pair, and an extended world, where rewards are defined per agent, consistent with the NR framework. In both cases, the axioms of efficiency, additivity, and null player uniquely characterize the corresponding Shapley analogs. But, these Shapley analogs are disjoint from the class of solutions satisfying the sequential axioms, even for convex and simple TCGs.
Our results reveal a fundamental tension in temporal cooperative games: when order matters, solution concepts satisfying natural sequential incentives must differ structurally from Shapley-based allocations, motivating new axiomatic foundations for sequential reward sharing.
1 Introduction
Organizations and institutions benefit from the complementary skill sets of their employees. However, the growth and overall value of an institution critically depend on the order in which employees join. Early joiners establish the vision and direction of the institution, which subsequently guides the hiring of core domain experts, followed by process enablers. Support staff and marketing personnel are then integrated in a sequence that aligns with the institution’s structure. A different order of joining can result in markedly different growth trajectories and institutional value. Classical cooperative games (Maschler et al., 2020) often overlook this phenomenon, as their worth functions are defined only for coalitions, ignoring the order of the agents within the coalition. Early works by Nowak and Radzik (1994); Sánchez and Bergantiños (1997) incorporate agent order through generalized characteristic functions (GCF). A series of follow-up papers (Van den Brink et al., 2014; Feng et al., 2013; Yuan and Driessen, 2013; Li et al., 2024; Michalak et al., 2014; Zou et al., 2020) consider extending or modifying some of their axioms and either characterize new solution concepts or provide alternative proofs. However, a primary feature of all these papers is that they consider solution concepts that are order-oblivious. The axiomatic characterization results try to extend the Shapley value for a player that does not depend on the order of arrival. In contrast, we focus on solution concepts that are order-sensitive, and our primary objective is to study the desirable properties of the order-sensitive solution concepts as well as their simultaneous satisfaction with Shapley-like properties that are order-sensitive. We will be referring to the cooperative games with generalized characteristic functions as temporal cooperative games (TCG) to distinguish them from classical cooperative games (CG) in the rest of the paper.
TCGs necessitate a new solution concept along with a corresponding set of desirable axioms. We argue that such a solution concept must be order-sensitive, since the same agent appearing in different positions of a sequence can induce different worths in TCGs, and therefore should receive different rewards. Motivated by this observation, we introduce three fundamental sequential properties, inspired by related notions in Ge et al. (2024). The first property, incentive for optimal arrival (I4OA), incentivizes agents to join in the optimal order that maximizes institutional worth.
The second property, online individual rationality (OIR), coincides with the notion introduced in Ge et al. (2024). OIR ensures that agents who join earlier receive weakly higher payoffs as additional agents arrive, thereby encouraging early participants to facilitate the arrival of others.
We also consider a standard third property, sequential efficiency (SE), which guarantees that the worth of any sequence is fully allocated among the agents in that sequence. This property is particularly important in settings where rewards must be distributed immediately upon agent arrivals, without waiting for future participants. In this paper, we characterize a class of solution concepts that satisfy these three properties. We further provide an approach illustrating how additional properties may restrict this class toward a unique solution concept.
Ideally, one would like such solution concepts to also satisfy Shapley-inspired order-sensitive axioms in TCGs. However, we prove that this is impossible: solution concepts that uniquely satisfy Shapley-inspired properties such as efficiency, additivity, and the null-player condition are disjoint from those satisfying TCG-appropriate properties such as I4OA, OIR, and efficiency (Figure˜3). Remarkably, this impossibility result persists even for restricted classes of TCGs, including convex and simple games, thereby underscoring a sharp departure from the classical cooperative game setting.
1.1 Our Contributions
The contributions of this paper are as follows:
  • We study cooperative games with generalized characteristic functions (GCF) introduced by Nowak and Radzik (1994), which we call temporal cooperative games (TCGs). We formalize axioms natural to this order-sensitive setting: incentive for optimal arrival (I4OA), online individual rationality (OIR), and sequential efficiency (SE) (Section˜2).
  • A key contribution is the introduction of sequential solution concepts (Section˜2). Unlike existing GCF literature, which yields order-oblivious reward shares to the players, our framework assigns rewards to every (sequence, player) pair, arguing that order-sensitive environments require order-sensitive solution concepts.
  • We characterize all solution concepts satisfying I4OA, OIR, and SE via a class of reward-sharing mechanisms SeqShare (Algorithm˜1), and identify a necessary condition, termed basis, for their satisfaction (Lemma˜1).
  • We formulate order-sensitive analogs of Shapley’s axioms for TCGs: efficiency, additivity, and null-player (Section˜4), and show that they uniquely characterize a Shapley-analogous sequential solution concept MargSol (Section˜5.1); symmetry follows as a consequence (Section˜6).
  • To connect with order-oblivious solutions, we define extended solution concepts by averaging over sequences. The extended axioms uniquely identify Ext-Shap (Section˜5.2), aligning exactly with the results in existing GCF literature.
  • Finally, we show that the extended solutions induced by SeqShare are disjoint from Ext-Shap, establishing a fundamental incompatibility between desirable sequential properties and Shapley-inspired axioms, even for convex and simple games (Section˜7).
1.2 Related Work
Classical cooperative game theory originated with the seminal work of Von Neumann and Morgenstern (1947). Soon thereafter, foundational solution concepts such as the core (Gillies, 1959), the nucleolus (Schmeidler, 1969), and the Shapley value (Shapley, 1953) were introduced. The study of online coalition formation is comparatively recent: Flammini et al. (2021) addressed the problem in its general form, Bullinger and Romen (2023) analyzed settings with random arrivals, and Bullinger and Romen (2025) examined stability in this context. In parallel, Lehrer and Scarsini (2013) studied dynamic cooperative games in which the worth of a coalition evolves over time, while Habis and Jean-Jacques Herings (2010); Kranich et al. (2005) extended the notion of the core to dynamic environments. Other strands of the literature focus on characterizing solution concepts that satisfy variants of the Shapley axioms for classical cooperative games. In particular, Van Den Brink (2002) employ efficiency, the null-player property, and fairness (appropriately defined for cooperative games) to characterize the Shapley value, while Casajus and Huettner (2013) modify the null-player property to obtain a broader class of solutions. Several other works propose alternative axiom systems characterizing the Shapley value (Hamiache, 2001; Casajus, 2014).
Online cooperative games, in which agents arrive sequentially and are incentivized to satisfy properties such as immediate participation, have been studied for monotone coalitional games by Ge et al. (2024) and for cost-sharing games by Zhang et al. (2025a). Axiomatic frameworks for online cooperative games were developed by Aziz et al. (2025), while Zhang et al. (2025b) proposed stable online coalition formation mechanisms aimed at maximizing social welfare. However, across all these models, worth is still defined over coalitions of agents rather than over sequences.
The notion of generalized characteristic functions defined on ordered subsets was first introduced by Nowak and Radzik (1994) and subsequently extended by Sánchez and Bergantiños (1997, 1999); Bergantiños and Sánchez (2001). Although these functions allow coalition worth to depend on order, the resulting solution concepts are sequence-independent. We refer to these as the NR and SB values, respectively. In particular, Nowak and Radzik (1994) extended Shapley-style axioms—such as additivity, the null-player condition, and efficiency—to temporal cooperative games (TCGs) and proposed a solution concept that uniquely satisfies them. Sánchez and Bergantiños (1997) introduced a weaker version of the null-player property than the one we consider (see Section˜4 for details) and required the additional assumption of symmetry to obtain a unique solution satisfying all four properties, which we do not impose.
In subsequent work on generalized characteristic functions, Yuan and Driessen (2013) characterize the SB value as the outcome of a bargaining game, while Feng et al. (2013) provide a characterization based on three properties, namely continuity, associated consistency, and the inessential game property. Van den Brink et al. (2014) introduce the order monotonicity axiom and characterize a class of solutions that includes the SB and NR values as extreme cases. Zou et al. (2020) and Li et al. (2024) study settings with restrictions on the order of agent arrivals and characterize solutions for these constrained generalized games. Finally, Michalak et al. (2014) implement generalized games as perfect-information games that admit the NR and SB values as subgame-perfect Nash equilibria.
The fundamental distinction between these works and ours lies in our focus on order-sensitive solution concepts. This perspective leads to two key departures: (1) the axioms of interest differ and are motivated by settings in which order-sensitive solutions are natural, and (2) the required properties are more stringent in the Shapley framework, as they must hold for every sequence $\pi$. Consequently, our main results are also different: (1) a characterization of order-sensitive axioms tailored to TCGs (Theorem˜1), (2) a characterization of Shapley-like yet order-sensitive axioms for TCGs (Theorem˜2), and (3) a demonstration of the disjointness of these two classes (Section˜7).
2 Preliminaries
Consider a set of agents $N=\{1,\ldots,n\}$. Denote $2^{N}$ to be the set of all possible subsets of $N$. Let $\textsc{perm}(S)$ denote all possible permutations of the players in $S$, where $S\in 2^{N}\setminus\emptyset$. Define all possible sequences of any length by $\Pi:=\{\pi:\pi\in\textsc{perm}(S),S\in 2^{N}\setminus\emptyset\}$. We denote the set of agents in a sequence $\pi$ by $P(\pi)$ and the player at position $i$ in $\pi$ as $\pi(i)$. We define a characteristic function $v:\Pi\to\mathbb{R}$, which assigns a worth to every sequence of agents. A temporal cooperative game (TCG) is thus described by a tuple $\langle N,v\rangle$. We also define $\Pi_{-N^{\prime}}:=\{\pi:\pi\in\textsc{perm}(S),S\in 2^{N\setminus N^{\prime}}\}$, i.e., the set of all agent sequences except the agents in set $N^{\prime}$. Also, denote the last player of a sequence $\pi$ as $\ell(\pi)$.
We call $\pi^{\prime}$ to be a prefix of $\pi$ if the first $|P(\pi^{\prime})|$ players of $\pi$ are the same as that of $\pi^{\prime}$, i.e., $\pi(i)=\pi^{\prime}(i),\forall i=1,2,\ldots,|P(\pi^{\prime})|$. A prefix $\pi^{\prime}$ of $\pi$ is denoted as $\pi^{\prime}\sqsubset\pi$. Note that $|P(\pi^{\prime})|<|P(\pi)|$, otherwise they become the same sequence. We assume that the characteristic function is monotone, i.e., $v(\pi^{\prime})\leqslant v(\pi)$ for all $\pi^{\prime}\sqsubset\pi$. We represent the predecessor of $i$ in $\pi$ with $\pi_{i}$, which denotes the longest prefix of $\pi$ not containing $i$. A sequence $\pi$ is a full sequence if it contains all the players, $P(\pi)=N$. Denote the set of all monotone characteristic functions by $V$.
The focus of this paper is to find desirable axioms under this setting and obtain an order-sensitive solution concept $\phi_{i},i\in N$, which divides the worth among the players for every sequence and satisfies these axioms. The term $\phi_{i}(\pi,v)$ denotes the share of value to agent $i\in P(\pi)$ when arriving in the sequence $\pi$ in the TCG $\langle N,v\rangle$. We denote by $\phi(\pi,v)$ the vector $(\phi_{i}(\pi,v),i\in N)$, with the convention that the solution concept assigns zero reward to the agents outside $\pi$, i.e., $\phi_{j}(\pi,v)=0,\ \forall j\in N\setminus P(\pi)$. For the rest of the paper, we use $\phi(\pi)$ to denote the reward vector, dropping $v$ when it is clear from context. We denote the optimal sequence $\pi^{*}(v)$ to be the one that maximizes the worth of this game, i.e., $\pi^{*}(v)\in\operatorname*{\arg\!\max}_{\pi\in\Pi}v(\pi)$. WLOG, we assume that the tie-breaking rule is such that $P(\pi^{*}(v))=N$ (with arbitrary tie-breaking among the full-length sequences), since the game is monotone. However, our results hold true even when $P(\pi^{*}(v))\subset N$, and the proof can be easily adapted.
We now define certain desirable properties specific to temporal cooperative games. The players arrive sequentially in a TCG, thus the rewards must also be given as they arrive. To an existing sequence $\pi$, when a new player $i$ arrives, we denote the new sequence by $\pi+i$. The worth generated by the sequence $\pi+i$ must be distributed among the players present in the sequence. Our first property ensures that the earlier players are incentivized when a new player joins, i.e., the solution concept should never decrease the reward of a player as new players join.
Definition 1 (Online Individual Rationality).
A solution concept $\phi$ is online individually rational (OIR) if for every TCG $\langle N,v\rangle$ $\phi_{i}(\pi,v)\leqslant\phi_{i}(\pi^{\prime},v),\forall i\in P(\pi),\forall\pi\sqsubset\pi^{\prime},\text{ where }\pi,\pi^{\prime}\in\Pi.$
The sequential nature of the game also creates the possibility of the players to stop arriving. Also, the next player’s arrival is uncertain beforehand. Therefore, the solution concept should ensure the worth of a given sequence $\pi$ is given out to all the players in $P(\pi)$ for every sequence $\pi$.
Definition 2 (Sequential Efficiency (SE)).
A solution concept $\phi$ is sequentially efficient if for every TCG $\langle N,v\rangle$ $\sum_{i\in P(\pi)}\phi_{i}(\pi,v)=v(\pi),\forall\pi\in\Pi.$
Since the number of agents is finite, every TCG will have an optimal sequence that generates the most value. We denote the optimal sequence of a TCG $\langle N,v\rangle$ by $\pi^{*}(v)$. Our next property incentivizes players to arrive in this optimal order. We use $\pi^{*}$ to denote the optimal sequence when the game $\langle N,v\rangle$ is clear from the context.
Definition 3 (Incentive for Optimal Arrival (I4OA)).
A solution concept $\phi$ satisfies incentive for optimal arrival (I4OA) if for every TCG $\langle N,v\rangle$ $\phi_{i}(\pi^{*}(v),v)\geqslant\phi_{i}(\pi,v),\forall\pi\in\Pi,i\in N.$
In the first part of this paper, we characterize reward-sharing mechanisms that satisfy OIR, I4OA, and SE for TCGs. In the second part, we define properties analogous to the celebrated Shapley properties in classical cooperative games and characterize the solution concepts that satisfy them.
3 The SeqShare Class of Mechanisms
In this section, we introduce a class of reward-sharing mechanisms that we later show to uniquely satisfy OIR, I4OA, and SE. As a precursor to this class, we first define two conditions called the basis conditions for a TCG $\langle N,v\rangle$.
Definition 4 (Basis conditions).
A TCG $\langle N,v\rangle$ satisfies basis conditions if there exists a vector $x\in\mathbb{R}^{n}$ such that the following conditions hold
$\begin{split}\textstyle(i)\sum_{i\in P(\pi)}x_{i}\geqslant v(\pi),\forall\pi\in\Pi,\ \ (ii)\sum_{i\in P(\pi^{*}(v))}x_{i}=v(\pi^{*}(v)).\end{split}$ (1)
We call the vector $x\in\mathbb{R}^{n}$ that satisfies the above conditions a basis solution.
The importance of the basis condition can be understood from the following result that shows that this condition is necessary for two desirable properties of TCGs. The basis conditions are analogous to the core of classical cooperative games. Certain desirable properties are contingent on these sets being non-empty, while it has a chance of being empty too.
Lemma 1.
For a TCG $\langle N,v\rangle$, if a solution concept $\phi$ satisfies I4OA and SE then there must be a basis solution.
Proof Sketch.  Given a TCG $\langle N,v\rangle$ and a solution concept $\phi$ satisfying I4OA and SE, choosing $x=\phi(\pi^{*})$ meets the basis conditions. Condition (ii) follows directly from SE applied to $\pi^{*}$. Condition (i) follows from I4OA, which bounds $\phi_{i}(\pi)$ by $\phi_{i}(\pi^{*})$ for all $i$ and $\pi$, and SE ensures $v(\pi)=\sum_{i}\phi_{i}(\pi)$, implying the required inequality. $\square$
Remark. The result above does not guarantee the existence of a solution concept satisfying I4OA and SE if a basis solution exists. In the following example, we show examples of solutions that satisfy the basis and use the basis solution as the reward for $\pi^{*}$, but still fall short of at least one of the given properties. Existence of a basis solution does not guarantee that trivial solution concepts can satisfy the $3$ properties together, as we illustrate in the following example. This shows that the method we propose needs non-trivial effort to ensure all three.
Example 1.
Consider the TCG: $v(1)=1,v(12)=2,v(2)=1,v(21)=3$. It is easy to check that the game has a basis solution. In this example, $\pi^{*}=21$, $x_{1}=1,x_{2}=2$. Now, consider the following solution concepts.
$\pi$ $v(\pi)$ $\phi$ $\phi^{\prime}$ $\phi^{\prime\prime}$
1 1 $(0,0)$ $(1,0)$ $(1,0)$
2 1 $(0,0)$ $(0,1)$ $(0,1)$
12 2 $(0,0)$ $(0,2)$ $(2,0)$
21 3 $(1,2)$ $(1,2)$ $(1,2)$
$\phi$ satisfies OIR and I4OA but not SE since $v(12)\neq\phi_{1}(12)+\phi_{2}(12)$. $\phi^{\prime}$ satisfies SE and I4OA but not OIR, $\phi(1)>\phi(12)$. $\phi^{\prime\prime}$ satisfies SE and OIR but not I4OA $\phi_{1}(12)>\phi_{1}(21)$. ∎
Our next endeavor is to construct a solution concept that will satisfy all of them.
Our desired class of solution concepts starts with the basis conditions to check if a solution exists. If the basis conditions hold, Algorithm˜1 yields a class of solution concepts that satisfy OIR, I4OA, and SE. It is a class because there are multiple possible solution concepts $\phi$ that can be given by this algorithm. However, we will show that each of them satisfies the three above desirable properties, and any solution concept satisfying the three is a valid solution concept given by this algorithm. The class SeqShare performs the following steps:
  1. 1.
    It checks and returns a basis solution if it exists. The returned value $x_{i}$ is treated as the upper bound for agent $i$’s reward for all sequences. Hence, $\phi_{i}(\pi^{*})$ is assigned $x_{i}$.
  2. 2.
    Given a sequence $\pi$ where $\pi(1)=i$, the algorithm assigns $\phi_{i}(i)=v(i)$. This value now serves as a lower bound for $\phi_{i}(\pi)$ for every sequence having $i$ as a prefix.
  3. 3.
    It computes the marginal contribution made by $\pi(2)=j$, and divides it among $i$ and $j$ such that their rewards $\phi_{i}(ij)$ and $\phi_{j}(ij)$ lie within the upper and lower bounds set for them. The lower bounds for the agents $i\text{ and }j$ are then updated to $\phi_{i}(ij)$ and $\phi_{j}(ij)$ respectively for every sequence containing $ij$ as a prefix.
  4. 4.
    The algorithm continues this iterative process of assigning rewards within the lower and upper bounds at the arrival of each agent, followed by updating the lower bound to the newest assigned value.
Figure˜1 provides a visual representation of SeqShare. The non-uniqueness of the solution concept given by SeqShare has two distinct sources. First, the basis solution may not be unique as Equation˜1 may be satisfied by multiple $x$ vectors. Second, given a basis solution, there may be multiple functions that are suitable candidates for $\phi$ as the Improvize function returns a non-unique $y$ vector. Hence, SeqShare is a class based on these two freedoms of choice. Once a solution concept $\phi$ is obtained via SeqShare, it can be implemented in an online fashion. This is because $\phi$ gives a reward share for every player in every sequence $\pi\in\Pi$. Hence, whenever an agent appears in any sequence, $\phi$ has a reward share for that agent at that point of arrival.
Algorithm 1 Class of Mechanisms SeqShare
1:TCG $\langle N,v\rangle$
2:Find $\pi^{*}(v)=\operatorname*{\arg\!\max}_{\pi}v(\pi)$, $x=\textsc{Basis}(N,v,\pi^{*}(v))$
3:if $x$ = null then Output null, exit
4:else $\phi(\pi^{*}(v))=x$; $\phi(\pi)=\mathbf{0},\ \forall\pi\neq\pi^{*}(v)$
5:end if
6:for $i\in N$ do $\phi_{i}(\{i\})=v(\{i\})$
7:end for
8:for $k=2,\ldots,n$ do
9:  for $\pi\in\Pi,\ |\pi|=k-1$ do
10:    for $j\in N\setminus P(\pi)$ do
11:     $y=\textsc{Improvize}(j,\pi,\phi(\pi),\phi(\pi^{*}(v)))$
12:     $\phi_{i}(\pi+j)=\phi_{i}(\pi)+y_{i},\forall i\in P(\pi)$; $\phi_{j}(\pi+j)=y_{j}$
13:    end for
14:  end for
15:end for
16:Output: $\phi$
1:procedure $\textsc{Basis}(N,v,\pi^{*}(v))$
2:  if solution to Equation˜1 exists then return $x$ of Equation˜1 w.r.t. $\pi^{*}(v)$
3:  else return null
4:  end if
5:end procedure
1:procedure $\textsc{Improvize}(j,\pi,\phi(\pi),\phi(\pi^{*}))$
2:  $y_{i}\in[0,\phi_{i}(\pi^{*})-\phi_{i}(\pi)],\ \forall i\in P(\pi+j)$
3:  $\sum_{i\in P(\pi+j)}y_{i}=v(\pi+i)-v(\pi)$
4:  return $y$
5:end procedure
P1P2P3P4Possible additional reward
P1P2P3P4Marginal Contribution of new player
P1P2P3P4
P1P2P3P4Reward already assigned
Figure 1: (a) The first player arrives, and all the value generated is assigned to player P1. (b) The second player P2 arrives, and the marginal value generated by the arrival of this player is distributed among P1 and P2 in a way that ensures the total reward of each agent is less than the upper bound. The total reward for any agent is the sum of the reward in the previous round and the share of the marginal increase assigned in the current round. (c) The third player P3 arrives, and the marginal value generated by P3 is distributed among P1, P2, and P,3 keeping total rewards lower than the upper bound. (d) Finally, P4 arrives, and the marginal value generated by this player can only be divided among P1,P2, and P4 as P3 has already reached its maximum possible reward.
Remark. While we use a basis solution as part of the algorithm, it should be noted that the existence of a basis solution is a property of a given TCG $\langle N,v\rangle$ and not that of a solution concept.
3.1 Properties of SeqShare
In this section, we present the results of the properties of SeqShare. We first show that the existence of a basis solution is sufficient to ensure a solution that satisfies OIR, I4OA, and SE using the procedures of SeqShare. Further, we show that if a game has a solution that satisfies OIR, I4OA, and SE, it is a solution concept in the SeqShare class. Lemma˜1 states that the existence of a basis solution is necessary to satisfy I4OA and SE. The following theorem completes this lemma with the OIR property.
Theorem 1.
Given a TCG $\langle N,v\rangle$, a reward function $\phi$ satisfies OIR, I4OA, and SE iff it is from the class SeqShare.
Proof Sketch.  If the game $\langle N,v\rangle$ violates the basis conditions, then by Lemma˜1 no solution concept can satisfy I4OA and SE, and SeqShare returns null. Hence, assume the basis conditions hold.
We first prove the reverse direction constructively using Algorithm˜1. The algorithm initializes the allocation using Basis on the maximal sequence $\pi^{*}(v)$ and assigns zero elsewhere. The construction proceeds by induction on sequence length.
For sequences of length one, SE holds trivially, I4OA follows from the basis conditions, and OIR holds by monotonicity. Assume the properties hold for all sequences of length k1k-1. For a sequence $\pi+i$ of length $k$, the function Improvize allocates the marginal reward so that payoffs are non-decreasing (OIR), bounded above by $\phi(\pi^{)}$ (I4OA), and sum to $v(\pi+i)$ (SE). Thus, the properties hold for all sequences.
For the forward direction, suppose there exists a solution concept $\phi$ satisfying I4OA, SE, and OIR but $\phi\notin\textsc{SeqShare}{}$. For all $\phi^{\prime}\in\textsc{SeqShare}{}$, let $\pi^{\prime}$ be the shortest sequence where $\phi^{\prime}$ and $\phi$ differ, among these choose the $\phi^{\prime}$ that corresponds to the longest such sequence. Let $\pi_{\textup{longshort}}$ be a shortest sequence where this $\phi^{\prime}$ and $\phi$ differ.
Using the hypothesis that $\phi$ satisfies OIR, I4OA, and Efficiency, we can construct a solution in SeqShare that matches $\phi$ on all sequences up to and including $|\pi_{\textup{longshort}}|$, contradicting the choice of $\pi_{\textup{longshort}}$. Hence, every solution concept satisfying I4OA, SE, and OIR belongs to SeqShare. $\square$
To position this class of solution concepts with that of the classical cooperative games, we consider the analogs of the Shapley properties in the context of temporal cooperative games.
4 Properties Inspired by Shapley in TCGs
Classical cooperative games and the classic solution concept due to Shapley operates on the setting where the characteristic function $v$ depends on a coalition. In TCG, the function $v$ depends on a sequence of players. To compare the solution concepts with that of the Shapley properties for a TCG $\langle N,v\rangle$, we first need to distinguish the solution concepts where it depends on the sequence $\pi$ (we will call this solution concepts as we did so far in this paper) and the ones where it does not depend on the sequence (we will call this extended solution concepts in the rest of this paper). The space of solution concepts is thus given by $\Phi=\{\phi:\Pi\times V\rightarrow\mathbb{R}^{n}\}$, and the space of extended solution concepts is given by $\Psi=\{\psi:V\rightarrow\mathbb{R}^{n}\}$, where $V$ is the set of all monotone characteristic functions. The sets $\Phi$ and $\Psi$ are therefore the order-sensitive and order-oblivious classes of solution concepts, respectively. A solution concept, however, can be reduced to an extended solution concept by averaging it over all possible full sequences, i.e., $\pi:P(\pi)=N$, as follows.
Definition 5.
Given a TCG $\langle N,v\rangle$ and a solution concept $\phi\in\Phi$, the reduction of $\phi$ to an extended solution concept $\overline{\phi}\in\Psi$ is defined as $\overline{\phi}_{i}=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi).$
In parity with the Shapley properties, we will now define the additivity, efficiency, and null player in the context of both types of solution concepts. To distinguish them, the properties for the solution concept that depends on the sequence are called sequential properties, while those for the extended solution concept are called extended properties. We begin with the sequential properties. The sequential efficiency property is already defined in Definition˜2.
Definition 6 (Sequential Additivity (SA)).
A solution concept $\phi\in\Phi$ is sequentially additive if for every pair of TCGs $\langle N,v\rangle$ and $\langle N,u\rangle$, $\phi_{i}(\pi,u)+\phi_{i}(\pi,v)=\phi_{i}(\pi,u+v),\ \forall\pi\in\Pi,\forall i\in N.$
Definition 7 (Sequential Null Player (SNP)).
A solution concept $\phi\in\Phi$ satisfies sequential null player if for every TCG $\langle N,v\rangle$ $v(\pi+i)=v(\pi),\ \forall\pi\in\Pi_{-i}\implies\phi_{i}(\pi)=0,\ \forall\pi\in\Pi.$
We now turn to the extended properties defined as follows.
Definition 8 (Extended Efficiency (EE)).
An extended solution concept $\psi\in\Psi$ satisfies extended efficiency if for every TCG $\langle N,v\rangle$ $\sum_{i\in N}\psi_{i}(v)=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi).$
The null player property in the extended space ensures that the player’s extended reward is zero.
Definition 9 (Extended Null Player (ENP)).
An extended solution concept $\psi\in\Psi$ satisfies extended null player if for every TCG $\langle N,v\rangle$ $v(\pi+i)=v(\pi),\ \forall\pi\in\Pi_{-i}\implies\psi_{i}(v)=0.$
Note that this definition of extended null player is stronger than that in (Sánchez and Bergantiños, 1997), where a player is null only if it does not improve the worth of a sequence if it is inserted in any position of the sequence, i.e., starting from the first, second, till the last position, while our null player only requires the agent to be added at the end. Since the ‘if’ condition of our definition is weaker, the property ENP is stronger than its null player property. Their characterization of the solution concepts requires four (and not three) properties, including symmetry, while we need only three, as we will soon see in Section˜5. We show in ˜2 that even if we use their definitions of the Shapley properties in this setting, the TCG-appropriate properties of solution concepts are in conflict with those of the Shapley-inspired properties, which is the fundamental conclusion of our paper.
Example 2.
Consider the TCG given by $v(1)=1,v(12)=1,(2)=1,v(21)=2$. The only solution that lies in SeqShare is $\phi(1)=(1,0),\phi(2)=(0,1),\phi(12)=(1,0),\phi(21)=(1,1)$. The reduction of this solution is $\overline{\phi}=(1,0.5)$. The solution characterized by Sánchez and Bergantiños (1997) for this game is $\psi=(0.75,0.75)$. This tells us that there exist games for which no $\phi\in\textsc{SeqShare}{}$ can be reduced to the solution given by Sánchez and Bergantiños (1997).
Additivity for an extended solution concept implies that the rewards must stay the same whether the agents play two games separately and consolidate their rewards or they play a game with the sum of the rewards of the two games.
Definition 10 (Extended Additivity (EA)).
An extended solution concept $\psi\in\Psi$ satisfies extended additivity if for every pair of TCGs $\langle N,v\rangle$ and $\langle N,u\rangle$ $\psi_{i}(u)+\psi_{i}(v)=\psi_{i}(u+v),\ \forall i\in N.$
We will drop the argument $v$ from $\psi(v)$ whenever it is clear from context.
One important property for Shapley value is symmetry. To define this property in the context of TCGs, we need the concept of swap. A swap of agents $i$ and $j$ in a sequence $\pi$ is: (a) $i$ replaced with $j$ if only $i\in P(\pi)$, or (b) $j$ replaced with $i$ if only $j\in P(\pi)$, or (c) both $i$ and $j$’s positions swapped if both $i,j\in P(\pi)$. This is denoted by $\pi_{\textup{swap}(i,j)}$.
Definition 11 (Sequential Symmetry (SS)).
A solution concept $\phi\in\Phi$ satisfies sequential symmetry if for every TCG $\langle N,v\rangle$, the following holds $\forall\pi\in\Pi$ $v(\pi)=v(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi\implies\phi_{i}(\pi)=\phi_{j}(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi.$
Definition 12 (Extended Symmetry (ES)).
An extended solution concept $\psi\in\Psi$ satisfies extended symmetry if for every TCG $\langle N,v\rangle$, $v(\pi)=v(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi\implies\psi_{i}=\psi_{j}.$
However, we will see in Section˜6 that each of these properties is a consequence of the other three properties in the two solution spaces.
Equipped with these definitions, we can now adapt the Shapley value to TCGs and draw connections to the solutions already discussed and those inspired by the idea of marginal contribution.
5 Solutions Satisfying Shapley-inspired Properties
In this section, we introduce two solution concepts, in $\Phi$ and $\Psi$ respectively. We show that both these solution concepts uniquely satisfy certain sets of desired properties defined in the previous section.
We define the solution concept $\textsc{MargSol}\in\Phi$ inspired by the notion of marginal contribution in classical cooperative games.
$\textsc{MargSol}_{i}(\pi)=v(\pi_{i}+i)-v(\pi_{i}),\ \forall\pi\in\Pi,\forall i\in N.$ (2)
Define the extended solution concept Extended Shapley Value by using the notion of marginal contribution in sequential cooperative games as follows.
$\begin{split}\textstyle\textsc{Ext-Shap}_{i}=&\textstyle\frac{1}{n!}\sum_{\pi\in\Pi:|\pi|=n}\textsc{MargSol}_{i}(\pi).\end{split}$ (3)
Observe that $\textsc{Ext-Shap}{}=\overline{\textsc{MargSol}}$, i.e., Ext-Shap is the reduced version of MargSol. The following lemma shows that the sequential properties are stronger than the extended properties, i.e., if a solution $\phi\in\Phi$ satisfies SE, SA, and SNP, its reduced solution $\overline{\phi}\in\Psi$ satisfies EE, EA, and ENP respectively.
Lemma 2.
For every TCG $\langle N,v\rangle$ and every solution concept $\phi\in\Phi$, the following implications hold.
  1. 1.
    $\phi$ satisfies SE $\implies\overline{\phi}$ satisfies EE,
  2. 2.
    $\phi$ satisfies SA $\implies\overline{\phi}$ satisfies EA, and
  3. 3.
    $\phi$ satisfies SNP $\implies\overline{\phi}$ satisfies ENP.
Proof Sketch.  Averaging the sequential property over all full-length sequences gives the desired extended property. $\square$
Our next goal is to show that the three properties are uniquely satisfied by MargSol. In order to show that, we use a construct similar to the Shapley uniqueness proof (Maschler et al., 2020, Chapter 19). We define the carrier games in the context of TCG as follows.
Definition 13 (Carrier Game).
For any sequence $\pi\in\Pi$, the carrier game over $\pi$ is the simple game [^1] We use the same terminology from the classical cooperative game: a simple game is a TCG where the worth of any sequence can either be zero or unity. $\langle N,u_{\pi}\rangle$ where:
$u_{\pi}(\pi^{\prime})=\begin{cases}1,&\text{if }\pi\sqsubset\pi^{\prime},\\ 0,&\text{otherwise}.\end{cases}$
We show that the carrier games span the space of all TCGs.
Lemma 3.
Every TCG $\langle N,v\rangle$ is a linear combination of carrier games.
Proof Sketch.  The proof follows similarly to the classical Shapley characterisation. The set of all carrier games forms a linearly independent set of size $|\Pi|$, thus forming a spanning set for the vector space of all TCGs. $\square$
5.1 Uniqueness of MargSol
To show that MargSol is the only solution in $\Phi$ that satisfies SE, SA, and SNP simultaneously, we need the following result.
Lemma 4.
Consider a TCG $\langle N,u_{\pi,\alpha}\rangle$ for a given $\pi\in\Pi$ and $\alpha$, where $u_{\pi,\alpha}$ is defined as follows.
$u_{\pi,\alpha}(\pi^{\prime})=\begin{cases}\alpha,&\text{if }\pi\sqsubset\pi^{\prime},\\ 0,&\text{otherwise}.\end{cases}$
If a solution concept $\phi\in\Phi$ satisfies SE and SNP, it must be true that
$\phi_{i}(\pi^{\prime},u_{\pi,\alpha})=\begin{cases}\alpha,&\text{if }i=\ell(\pi)\text{ and }\pi\sqsubset\pi^{\prime},\\ 0,&\text{otherwise}.\end{cases}$
Intuitively, this result says that all the players except the last player of $\pi$, $\ell(\pi)$, are null players in TCG $\langle N,u_{\pi,\alpha}\rangle$. Due to efficiency, it is necessary that the last player must get the whole reward of $\alpha$. The formal proof is as follows.
Proof.
Suppose $i\in N\setminus\{\ell(\pi)\}$ is not the last player of $\pi$. In $\langle N,u_{\pi,\alpha}\rangle$, every such player satisfies $u_{\pi,\alpha}(\pi^{\prime}+i)=u_{\pi,\alpha}(\pi^{\prime}),\ \forall\pi^{\prime}\in\Pi_{-i}$, i.e., each such $i$ is a sequential null player. Until player $\ell(\pi)$ arrives in the sequence $\pi$, no marginal contribution is generated in this TCG by any other player arriving after. Hence, by SNP, all other players must get $\phi_{i}(\pi^{\prime},u_{\pi,\alpha})=0,\forall\pi^{\prime}\in\Pi$. The entire worth of $\alpha$ then must go to $\ell(\pi)$ due to SE. ∎
Theorem 2.
For every TCG $\langle N,v\rangle$, a solution concept $\phi\in\Phi$ satisfies SE, SA, and SNP iff $\phi\equiv\textsc{MargSol}$.
Proof Sketch.  ($\Leftarrow$) If player $i$ is a sequential null player, then for all $\pi,\ \textsc{MargSol}_{i}(\pi)=0$. Additivity (SA) holds by linearity of marginal contributions, and efficiency (SE) follows from telescoping the sum of incremental values across the sequence.
($\Rightarrow$) Any TCG can be written as a sum of carrier games. For each carrier game, MargSol is the unique solution satisfying SE and SNP. If another solution $\phi$ also satisfies SE, SA, and SNP, it must agree with MargSol on all carrier games. By linearity and the basis decomposition, $\phi$ must equal MargSol on the entire game. $\square$
5.2 Uniqueness of Ext-Shap
Ext-Shap is an averaged version of MargSol (Equation˜3), and therefore the mapping from MargSol to Ext-Shap need not be unique. Consequently, although MargSol uniquely satisfies SE, SA, and SNP (Theorem˜2), this does not automatically imply that Ext-Shap uniquely satisfies EE, EA, and ENP. However, Nowak and Radzik (1994, Theorem 1) established that Ext-Shap indeed uniquely satisfies EE, EA, and ENP. We restate this theorem below and present an alternative proof in the appendix using the framework developed in this paper. Our characterization of order-sensitive solution concepts thus aligns with, and complements, the characterization of order-oblivious solution concepts existing in the literature.
Theorem 3 (Nowak and Radzik (1994)).
For every TCG $\langle N,v\rangle$, an extended solution concept $\psi\in\Psi$ satisfies EE, EA, and ENP iff $\psi=\textsc{Ext-Shap}{}$.
Remark. The Shapley value equivalents in the TCG world, MargSol and Ext-Shap, satisfy ‘almost’ similar uniqueness properties like the Shapley value. However, a careful reader can spot the difference since Shapley value was unique for four (and not three as we did here) properties, which also included symmetry. The TCG-equivalent definitions of symmetry are given in Definitions˜11 and 12. In Section˜6, we show why these two properties were not necessary to obtain the uniqueness theorems Theorems˜2 and 3.
6 Why Symmetry Is Not Necessary?
Properties in the TCG framework are more restrictive as they make claims for all sequences, making them stronger than claims for all sets, as in the classical cooperative games. Hence, three properties are sufficient to ensure unique solution concepts like MargSol and Ext-Shap. In the following, we show that MargSol and Ext-Shap satisfy SS (Definition˜11) and ES (Definition˜12) respectively, defined in the spirit of TCGs. These results show that the three other properties imply the symmetry property.
Theorem 4.
MargSol and Ext-Shap satisfy SS and ES respectively.
Proof.
Part 1 (MargSol): Consider a TCG $\langle N,v\rangle$, s.t. $v(\pi)=v(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi$. We show the following $\forall\pi\in\Pi,\forall i\in N$
$\textsc{MargSol}_{i}(\pi)$
$=v(\pi_{i}+i)-v(\pi_{i})$
$=v((\pi_{i})_{\textup{swap}(i,j)}+j)-v((\pi_{i})_{\textup{swap}(i,j)})$
$=v((\pi_{\textup{swap}(i,j)})_{j}+j)-v((\pi_{\textup{swap}(i,j)})_{j})$
$=\textsc{MargSol}_{j}(\pi_{\textup{swap}(i,j)}).$
The first equality follows from definition. The second equality comes from the ‘if’ condition of the SS property. The third equality follows from the fact that $(\pi_{i})_{\textup{swap}(i,j)}=(\pi_{\textup{swap}(i,j)})_{j}$. The last equality is again by definition. Hence, we show that MargSol satisfies SS.
Part 2 (Ext-Shap): Since $\textsc{Ext-Shap}{}_{i}=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\textsc{MargSol}_{i}(\pi)$, and we have already shown in Part 1 that $\textsc{MargSol}_{i}(\pi)=\textsc{MargSol}_{j}(\pi_{\textup{swap}(i,j)})$, it is straightforward to show that
$\textsc{Ext-Shap}{}_{i}$
$=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\textsc{MargSol}_{i}(\pi)$
$=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\textsc{MargSol}_{j}(\pi_{\textup{swap}(i,j)})$
$=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\textsc{MargSol}_{j}(\pi)$
$=\textsc{Ext-Shap}{}_{j}$.
The third equality holds since $\pi_{\textup{swap}(i,j)}$ just changes the order of summation in the set $\{\pi\in\Pi:P(\pi)=N\}$. Hence, Ext-Shap satisfies ES. ∎
Remark. In classical cooperative games, the Shapley value is the only solution concept satisfying efficiency, additivity, null player property and symmetry. Symmetry is required to make the solution concept unique. The difference in our setting lies in the definition of the null player property for TCGs.
In the classical setting, a carrier game is defined with respect to a critical set of players. When the set is present, only then does a coalition generate unit positive worth. Dividing the worth among the non-null players requires a property like symmetry.
In TCG framework, a carrier game corresponds to a critical sequence of players. When this critical sequence happens to be a prefix of a given sequence, only then can it generate a unit positive worth. As a result, only the last player of the critical sequence is non-null, so symmetry is not required. Only one player gets all the worth.
7 SeqShare and Shapley-compliant: Best of Both Worlds
The SeqShare class characterizes three most desirable properties for TCGs: OIR, SE, and I4OA. On the other hand, Ext-Shap uniquely characterizes the three most desirable Shapley-equivalent properties: EA, EE, and ENP. A solution concept that satisfies the temporal properties of TCGs and simultaneously has its reduced version satisfying the Shapley properties would therefore be the ideal candidate in a TCG. In this section, we show that such a “best of both worlds” solution concept does not exist, even in special games like convex or simple games (Maschler et al., 2020, Chapter 19).
In order to show the result for general TCGs, we first make the following observation about solution concepts in SeqShare that reduce to Ext-Shap.
Lemma 5.
For a TCG $\langle N,v\rangle$ and a solution concept $\phi\in\textsc{SeqShare}{}$, if $\overline{\phi}=\textsc{Ext-Shap}{}$ then the following holds $\forall i\in N$
$\phi_{i}(\pi^{*}(v))\geqslant\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi_{i}+i)-v(\pi_{i})=\textsc{Ext-Shap}{}_{i}$, (4)
where $\pi^{*}(v)$ is the optimal sequence of the TCG.
Proof Sketch.  Assume $\phi\in\textsc{SeqShare}$ with $\overline{\phi}=\textsc{Ext-Shap}$. Suppose $\phi_{i}(\pi^{*}(v))$ is strictly less than the average marginal contribution. By I4OA, $\phi_{i}(\pi)\leq\phi_{i}(\pi^{*}(v))$ for all $\pi$, so the average $\overline{\phi}_{i}$ is also strictly less than $\textsc{Ext-Shap}_{i}$. This contradicts the assumption $\overline{\phi}=\textsc{Ext-Shap}$, hence no such $\phi$ can exist. $\square$
Hence, we conclude that Equation˜4 is necessary to ensure that a solution from SeqShare reduces to Ext-Shap. In the example given below, we show a game that does not satisfy Equation˜4, and therefore, there does not exist any solution concept $\phi\in\textsc{SeqShare}{}$ that has a reduction to Ext-Shap.
Example 3.
123v(1)=1v(1)=1v(12)=3v(12)=3v(123)=av(123)=a
132v(1)=1v(1)=1v(13)=4v(13)=4v(132)=6v(132)=6
213v(2)=bv(2)=bv(21)=4v(21)=4v(213)=8v(213)=8
231v(2)=bv(2)=bv(23)=5v(23)=5v(231)=7v(231)=7
312v(3)=4v(3)=4v(31)=5v(31)=5v(312)=av(312)=a
321v(3)=4v(3)=4v(32)=5v(32)=5v(321)=7v(321)=7
Figure 2: A generic TCG serving as counterexample for multiple settings for different values of $a$ and $b$, e.g., no solution in SeqShare can be reduced to Ext-Shap.
Consider the TCG shown in Figure˜2 with $a=7,b=3$. Clearly, $v(\pi^{*})=8$. The basis solution (Equation˜1) dictates that $x_{1}\geqslant v(1)=1,x_{2}\geqslant v(2)=3,x_{3}\geqslant v(3)=4$ and $x_{1}+x_{2}+x_{3}=8$, which implies that the basis solution must be $x_{1}=1,x_{2}=3,x_{3}=4$. This is the only possible basis solution and therefore any $\phi\in\textsc{SeqShare}{}$ assigns $\phi_{i}(\pi^{*})=x_{i}$. Routine calculations for Player 1 gives $\textsc{Ext-Shap}{}_{1}=\frac{8}{6}$. However, $\phi_{1}(\pi^{*})=1<\frac{8}{6}=\textsc{Ext-Shap}{}_{1}$, which violates Equation˜4. ∎
Remark. Since no solution in SeqShare can be reduced to Ext-Shap, no solution in SeqShare can be MargSol. In particular, MargSol satisfies SE by construction. Since we are considering only monotone games, the marginal contribution, $v(\pi_{i}+i)-v(\pi_{i})$, is always non-negative for every player $i\in N$, and remains so even after the arrival of future agents, which ensures OIR. Hence, the only property that MargSol violates is I4OA. The following example illustrates this. This rules out the possibility of obtaining a solution concept that satisfies both sets of properties in both the original and extended spaces.
[…]
(a) solutions in the extended space
SNP $\cap$ SAMargSolSEI4OAOIRSeqSh
(b) solutions in the sequential space
Figure 3: Impossibilities of Shapley-inspired properties and solutions in SeqShare(SeqSh).
Since the Shapley-inspired properties and I4OA are incompatible for general games, we aim to find an intersection in the following sections in special classes of TCGs.
7.1 Convex Games
In classical cooperative games, the Shapley value exhibits some nice properties in the class of convex games. One such property is that for convex games, the Shapley Value always lies in the core. Since the basis conditions (Equation˜1) are analogous to the core conditions in classical games, we check for a similar result in the case of TCGs. We also check if there is a solution in SeqShare that reduces to Ext-Shap in this class of games. For that, first we need to define convex TCGs.
Definition 14 (Convex Temporal Cooperative Games).
A TCG $\langle N,v\rangle$ is convex if the following holds $v(\pi+i)-v(\pi)\leqslant v(\pi^{\prime}+i)-v(\pi^{\prime}),\ \forall\pi^{\prime},\pi\text{ s.t. }i\notin P(\pi^{\prime})\text{ and }P(\pi)\subseteq P(\pi^{\prime}).$
The answer to the question of having a solution in SeqShare that reduces to Ext-Shap is unfortunately negative. First, the following example illustrates a convex TCG that has no basis solution, unlike the convex games in classical cooperative game where it is always in the core.
Example 4.
Consider the convex TCG as shown in ˜3 with $a=7,b=4$. The basis solution (Equation˜1) must satisfy $x_{1}\geqslant 1,x_{2}\geqslant 4,x_{3}\geqslant 4$ and $x_{1}+x_{2}+x_{3}=8$, that are clearly impossible. Hence, the given convex TCG has no basis solution. ∎
Our next example shows that even for a convex TCG having a basis solution, there may be no solution in SeqShare that reduces to Ext-Shap.
Example 5.
Consider the TCG as used in ˜3, the game is convex and violates Equation˜4. ∎
Due to space constraints, we define simple TCGs and explore the connections between Ext-Shap and SeqShare in Section˜7.2.
7.2 Simple Games
In classical cooperative games, simple games or ‘0-1’ games are defined as the games where the worth of a coalition can either be 0 or $1$. In TCGs, we follow an analogous definition where the worth of any sequence can either take 0 or $1$ values. We define simple TCGs as follows.
Definition 15 (Simple Temporal Cooperative Games).
A TCG $\langle N,v\rangle$ with $v:\Pi\rightarrow\{0,1\}$ is called a simple temporal cooperative game.
Since we consider monotone games in this paper, monotonicity for simple TCGs implies that for all $\forall\pi,\pi^{\prime}\in\Pi$ such that $\pi\sqsubset\pi^{\prime}$, it holds that $v(\pi)=1\implies v(\pi^{\prime})=1$.
The following example shows that even for simple games, there may be no solution in SeqShare that reduces to Ext-Shap.
Example 6.
Consider the simple TCG given in Figure˜4. The unique basis solution (Equation˜1) is $x_{1}=1,x_{2}=0,x_{3}=0$. Therefore any $\phi\in\textsc{SeqShare}{}$ assigns $\phi_{i}(\pi^{*})=x_{i}$. Routine calculations for Player 2 gives $\textsc{Ext-Shap}{}_{2}=\frac{1}{6}$. However, $\phi_{2}(\pi^{*})=0<\frac{1}{6}=\textsc{Ext-Shap}{}_{2}$, which violates Equation˜4. ∎
123v(1)=1v(1)=1v(12)=1v(12)=1v(123)=1v(123)=1
132v(1)=1v(1)=1v(13)=1v(13)=1v(132)=1v(132)=1
213v(2)=0v(2)=0v(21)=0v(21)=0v(213)=1v(213)=1
231v(2)=0v(2)=0v(23)=1v(23)=1v(231)=1v(231)=1
312v(3)=0v(3)=0v(31)=0v(31)=0v(312)=1v(312)=1
321v(3)=0v(3)=0v(32)=0v(32)=0v(321)=1v(321)=1
Figure 4: A simple TCG for which no solution in SeqShare reduces to Ext-Shap.
Even for simple TCGs, we cannot hope to have a unique basis solution nor a solution in SeqShare that reduces to Ext-Shap.
8 Conclusions and Future Work
TCGs bring in new challenges for the axioms and depart from classical cooperative games. The solution concepts proposed in this paper were order-sensitive. We characterized three natural temporal properties, namely incentive for optimal arrival (I4OA), online individual rationality (OIR), and sequential efficiency (SE), with a class of reward-sharing mechanisms. The Shapley analog for order-sensitive solution concepts uniquely characterizes efficiency, additivity, and the null player property. Importantly, we demonstrated a fundamental incompatibility: the Shapley analogs and the mechanisms satisfying I4OA, OIR, and SE are disjoint, and this conflict persists even in restricted settings such as convex and simple TCGs.
Our results highlight a key tension in extending classical cooperative game theory to generalized characteristic functions: desirable temporal properties cannot be reconciled with Shapley-like axioms. As future work, an important direction is to investigate whether a unique solution concept can be obtained by imposing additional properties beyond I4OA, OIR, and SE. Another avenue is to identify special classes of temporal cooperative games where Shapley-like properties and temporal properties such as I4OA can coexist. These questions open new opportunities for developing principled solution concepts for cooperative games in sequential environments.
Appendix A Appendix to Section˜3
A.1 Proof of Lemma˜1
Proof.
Consider a TCG $\langle N,v\rangle$ and a solution concept $\phi$ that satisfies I4OA and SE in the game. We claim that $x=\phi(\pi^{*})$ satisfies the basis condition.
  • Condition (ii) of Equation˜1 is satisfied by applying SE of $\phi$ for $\pi^{*}$.
  • Since $\phi$ also satisfies I4OA, $\phi_{i}(\pi)\leqslant\phi_{i}(\pi^{*}),\forall\pi\in\Pi,\forall i\in N$. Hence, $\sum_{i\in P(\pi)}\phi_{i}(\pi)\leqslant\sum_{i\in P(\pi)}\phi_{i}(\pi^{*})$. By SE of $\phi$, $v(\pi)=\sum_{i\in P(\pi)}\phi_{i}(\pi)$. Therefore, we get $\sum_{i\in P(\pi)}\phi_{i}(\pi^{*})\geqslant v(\pi)$, which is condition (i) of Equation˜1.
Hence $\phi(\pi^{*})$ satisfies basis conditions. ∎
A.2 proof of Theorem˜1
Proof.
If the game $\langle N,v\rangle$ does not satisfy the basis conditions, then from Lemma˜1 we know that no $\phi$ can satisfy I4OA and SE. In that case, SeqShare returns null and hence the theorem holds. For the rest of the proof, we will assume that the game satisfies the basis conditions.
First, we prove the reverse direction, constructively via Algorithm˜1. The function Basis returns a valid partial allocation $x$, which is assigned to the value of the solution concept for the specific sequence $\pi^{*}(v)$, i.e., $\phi(\pi^{*}(v))$, and $\phi$ is initialized to zero for all other sequences (including the empty sequence). The rest of the proof is via induction on the length of the sequence $\pi$.
First, observe that for all unit-length sequences, the allocation given by Algorithm˜1 trivially satisfies SE. Property I4OA is satisfied due to the fact that $\phi_{i}(\{i\})=v(\{i\})\leqslant\phi_{i}(\pi^{*}(v)),\forall i\in N$, given by the Basis conditions. OIR is trivial too, since the game is monotone and $\phi_{i}(\emptyset)=0,\forall i\in N$.
Hence, by the induction hypothesis, assume that the forward implication of the theorem is true for all $\pi$ such that $|\pi|=k-1,k\geqslant 2$. We show that the implication is true for every $\pi+i$ of length $k$ as well. Consider the function Improvize. For every $\pi$ of length k1k-1 and the existing partial share $\phi(\pi)$, Improvize allocates the reward to agents in such a way that it is at least $\phi_{i}(\pi)$ (required by OIR) and at most $\phi_{i}(\pi^{*})$ (required by I4OA). This allocation is always possible since the monotonicity of the game ensures $v(\pi+i)\geqslant v(\pi)$ and $v(\pi+i)\leqslant v(\pi^{*})$ (by definition of $\pi^{*}$). Hence, I4OA and OIR are satisfied for each $i\in N$. Also, the construction of the $y$ vectors (reward shares) is such that the sum of rewards of all agents $j\in\pi+i$ equals $v(\pi+i)$, satisfying SE.
Next, we prove the forward direction. We prove this via contradiction. Suppose there is a solution concept $\phi$ for a given game $\langle N,v\rangle$ that satisfies I4OA, SE, and OIR but is not from the class SeqShare. Then we will construct a solution concept $\tilde{\phi}\in\textsc{SeqShare}{}$ which matches $\phi$ for every sequence and for every player, leading to the contradiction that $\phi\in\textsc{SeqShare}{}$.
For every solution $\phi^{\prime}$ from class SeqShare, there is a shortest length sequence $\pi\in\Pi$ such that $\phi^{\prime}_{i}(\pi)\neq\phi_{i}(\pi)$ for some $i\in P(\pi)$ while $\phi^{\prime}_{i}(\pi^{\prime})=\phi_{i}(\pi^{\prime})$ for all $i\in P(\pi^{\prime})$ and $\pi^{\prime}\in\Pi:|\pi^{\prime}|<|\pi|$, i.e., $\pi$ has the minimum length where the solution concepts $\phi$ and $\phi^{\prime}$ differ.
Choose the $\phi^{\prime}$ where the length of the sequence $\pi$ is the longest over all $\phi^{\prime}$s (breaking ties arbitrarily). Call this longest (over solution concepts) shortest (over sequences) sequence to be $\pi_{\textup{longshort}}$. Note that, by assumption, no $\phi^{\prime\prime}\in\textsc{SeqShare}{}$ can satisfy $\phi^{\prime\prime}_{i}(\pi)=\phi_{i}(\pi)$ for all $i\in P(\pi)$ where $\pi\in\Pi$ and $|\pi|\leqslant|\pi_{\textup{longshort}}|$. We will construct a $\tilde{\phi}$ that achieves this and belongs to SeqShare to show a contradiction.
Start with a game where the required properties are satisfiable. Since this game has a solution concept $\phi$ that satisfies I4OA and SE, Basis must be non-empty (Lemma˜1). Also, $\phi(\pi^{*})$ will be a solution to the basis by Lemma˜1. Hence, set $\tilde{\phi}$ such that $\tilde{\phi}(\pi^{*})=\phi(\pi^{*})$.
Construct $\tilde{\phi}\in\textsc{SeqShare}{}$ as follows: $\tilde{\phi}_{i}(\pi^{\prime})=\phi_{i}(\pi^{\prime}),\forall i\in P(\pi^{\prime}),\forall\pi^{\prime}:|\pi^{\prime}|<|\pi_{\textup{longshort}}|$. Clearly, this is feasible, since there are $\phi^{\prime\prime}\in\textsc{SeqShare}{}$ that satisfy these conditions and the construction of $\phi^{\prime\prime}$ is always from shorter to longer sequences.
Since $\phi$, the solution concept not in SeqShare, satisfies OIR, $\phi_{i}(\pi_{\textup{longshort}})-\phi_{i}(\pi_{\textup{longshort}}-\ell(\pi_{\textup{longshort}}))=y_{i}\geqslant 0$ for all $i\in P(\pi_{\textup{longshort}})$, where $\ell(\pi)$ is the last agent of the sequence $\pi$ and the subtraction $\pi-\ell(\pi)$ denotes that the last agent of $\pi$ is removed from the end of that sequence. Since, $\phi$ satisfies I4OA, $\phi_{i}(\pi_{\textup{longshort}})\leqslant\phi_{i}(\pi^{*})$ for all $i\in P(\pi_{\textup{longshort}})$.
Now, set, $\forall i\in P(\pi_{\textup{longshort}})$
$\tilde{\phi}_{i}(\pi_{\textup{longshort}})$ $=\tilde{\phi}_{i}(\pi_{\textup{longshort}}-\ell(\pi_{\textup{longshort}}))+y_{i}$
$=\phi_{i}(\pi_{\textup{longshort}}-\ell(\pi_{\textup{longshort}}))+y_{i}$
$=\phi_{i}(\pi_{\textup{longshort}}).$
Note that this retains the solution concept $\tilde{\phi}$ in SeqShare since it adds $y_{i}$ on top of the previously allocated reward to the sequence $\pi_{\textup{longshort}}-\ell(\pi_{\textup{longshort}})$, since the value that $y_{i}$ can take is according to the function Improvize. After adding $y_{\ell(\pi_{\textup{longshort}})}$ it gives $\phi_{\ell(\pi_{\textup{longshort}})}(\pi_{\textup{longshort}})$ which is at most $\phi_{\ell(\pi_{\textup{longshort}})}(\pi^{*})=\tilde{\phi}_{\ell(\pi_{\textup{longshort}})}(\pi^{*})$. The second equality above comes because by construction $\tilde{\phi}$ matches $\phi$ for all sequences of length less than that of $\pi_{\textup{longshort}}$.
Since $\pi_{\textup{longshort}}$ was chosen arbitrarily among all possible longest shortest sequences, and this process of assigning $\tilde{\phi}_{i}(\pi_{\textup{longshort}})$ does not affect other sequences of the same length, we can do the same for all other sequences of the same length simultaneously. Hence, $\tilde{\phi}$ matches $\phi$ for all sequences of that length and lower. This completes the proof. ∎
Appendix B Appendix to Section˜5
B.1 Proof of Lemma˜2
Proof.
Part (a): since $\phi$ satisfies SE, we have $\sum_{i\in P(\pi)}\phi_{i}(\pi,v)=v(\pi),\forall\pi\in\Pi$. Averaging this over all full-length sequences, we get $\forall\pi\in\Pi$
$\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi)$ $=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\sum_{i\in P(\pi)}\phi_{i}(\pi,v)$
$=\sum_{i\in N}\left(\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi,v)\right)$
$=\sum_{i\in N}\overline{\phi}_{i}(\pi,v)$
The first equality comes via changing the order of the summation, and the second equality is from the definition of the reduced solution concept. Hence, $\overline{\phi}$ satisfies EE.
Part (b): since $\phi$ satisfies SA, we have $\phi_{i}(\pi,u)+\phi_{i}(\pi,w)=\phi_{i}(\pi,u+w),\forall i\in N$. Averaging this over all full length sequences, we get
$\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}(\phi_{i}(\pi,u)+\phi_{i}(\pi,w))$
$=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi,u+w),\>\forall\pi\in\Pi,i\in N$
$\implies\overline{\phi}_{i}(u)+\overline{\phi}_{i}(w)=\overline{\phi}_{i}(u+w),\forall i\in N$
The implication holds from the definition of the reduced solution concept. Hence, $\overline{\phi}$ satisfies EA.
Part (c): since $\phi$ satisfies SNP, we have $v(\pi+i)=v(\pi),\>\forall\pi\in\Pi:i\notin P(\pi)\implies\phi_{i}(\pi,v)=0,\forall\pi\in\Pi$. For every such player $i$, averaging over all possible full length sequences, we get $\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi,v)=0$, which is $\overline{\phi}_{i}(v)=0$ by definition. Hence,
$v(\pi+i)=v(\pi),\>\forall\pi\in\Pi:i\notin P(\pi)\implies\overline{\phi}_{i}(v)=0.$
Hence, $\overline{\phi}$ satisfies ENP. ∎
B.2 Proof of Lemma˜3
Proof.
Consider an arbitrary TCG $\langle N,v\rangle$. We show that every $v$ can be written as a linear combination of $u_{\pi}$ (Definition˜13).
Part 1: we show that $u_{\pi},\pi\in\Pi$ are linearly independent. Suppose not, then $\exists\{\alpha_{\pi}:\pi\in\Pi\}$, not all 0 such that $\sum_{\pi\in\Pi}\alpha_{\pi}u_{\pi}(\pi^{\prime})=0,\forall\pi^{\prime}\in\Pi$.
Let $T=\{\alpha_{\pi}:\pi\in\Pi,\alpha_{\pi}\neq 0\}$ be the set of all non-zero $\alpha_{\pi}$’s. Let $\tau=\{\pi:\pi\in\Pi,\alpha_{\pi}\in T\}$ be the indices of $\alpha$ in $T$. Observe that, any $\pi\notin\tau$ must have $\alpha_{\pi}=0$. Consider the permutation $\pi_{0}\in\tau$, such that no prefix of $\pi_{0}$ is in $\tau$. We call this a minimal sequence of $\tau$. We can write
$\sum_{\pi\in\Pi}\alpha_{\pi}u_{\pi}(\pi_{0})=$ $\sum_{\pi\in\Pi,\pi\sqsubset\pi_{0}}\alpha_{\pi}u_{\pi}(\pi_{0})+\alpha_{\pi_{0}}u_{\pi_{0}}(\pi_{0})$
$+\sum_{\pi\in\Pi,\pi\not\sqsubset\pi_{0}}\alpha_{\pi}u_{\pi}(\pi_{0})=0.$
[…]
Part 2: we next prove that $u_{\pi}$ span the entire space of $\mathbb{R}^{|\Pi|}$. Every sequence $\pi\in\Pi$ denotes a unique carrier game $\langle N,u_{\pi}\rangle$. Hence, the number of unique $u_{\pi}$ are $|\Pi|$. $u_{\pi}:\Pi\to\mathbb{R}$, so if they are linearly independent, they must span the entire vector space of $|\Pi|$ dimensions, i.e., $\mathbb{R}^{|\Pi|}$.
Note that from parts 1 and 2, the lemma is immediate, since for any arbitrary TCG $\langle N,v\rangle$, the worth function $v$ lives in $\mathbb{R}^{|\Pi|}$, since $v:\Pi\to\mathbb{R}$. For a fixed $N$, $\Pi$ is finite. Parts 1 and 2 showed that $u_{\pi},\pi\in\Pi$ forms a basis of $\mathbb{R}^{|\Pi|}$, and hence any $v$ can be written as a linear combination of $u_{\pi}$. Therefore, every TCG $\langle N,v\rangle$ is a linear combination of the carrier games. ∎
B.3 Proof of Theorem˜2
Proof.
($\Leftarrow$) It is trivial to see that if $i$ is a sequential null player, $\textsc{MargSol}_{i}(\pi)=0,\forall\pi\in\Pi$. Also, $\textsc{MargSol}_{i}(\pi,u)+\textsc{MargSol}_{i}(\pi,v)=u(\pi_{i}+i)+v(\pi_{i}+i)-u(\pi_{i})-v(\pi)=\textsc{MargSol}_{i}(\pi,u+v)$. Hence MargSol satisfies SA and SNP. To see that it also satisfies SE, note that
$\sum_{i\in P(\pi)}\textsc{MargSol}_{i}(\pi)=\sum_{i\in P(\pi)}\left(v(\pi_{i}+i)-v(\pi_{i})\right)$
$=(v(\pi(1))-v(\emptyset))+(v(\pi(1)\pi(2))-v(\pi(1)))$
$+(v(\pi(1)\pi(2)\pi(3))-v(\pi(1)\pi(2)))+$
$\ldots+(v(\pi(1)\pi(2)\pi(3)\cdots\pi(|P(\pi)|))$
$-v(\pi(1)\pi(2)\cdots\pi(|P(\pi)|-1)))=v(\pi).$
Where $\pi(k)$ is the agent at the $k^{\textup{th}}$ position of $\pi$.
($\Rightarrow$) Consider an arbitrary TCG $\langle N,v\rangle$. Write it as a sum of its carrier games which is possible due to Lemma˜3:
$v(\pi^{\prime})=\sum_{\pi\in\Pi}\alpha_{\pi}u_{\pi}(\pi^{\prime})=\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}(\pi^{\prime}).$ (5)
For each carrier game, we know the structure of the unique solution concept that satisfies SE and SNP from Lemma˜4. Note that MargSol satisfies all three properties for every game $\langle N,v\rangle$, hence it must satisfy them for the carrier games as well. Suppose there exists a different solution concept $\phi$ that also satisfies the same three properties. We will prove that $\phi$ must be MargSol.
From Lemma˜4 and the discussion above, we conclude that $\phi$ must be the same as MargSol for the carrier game $\langle N,u_{\pi,\alpha}\rangle$. Hence we have $\forall i\in N,\forall\pi,\pi^{\prime}\in\Pi$
$\textsc{MargSol}_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}})=\phi_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}}),$
$\Rightarrow\sum_{\pi\in\Pi}\textsc{MargSol}_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}})=\sum_{\pi\in\Pi}\phi_{i}(\pi^{\prime},u_{\pi,\alpha_{\pi}}),$
$\Rightarrow\textsc{MargSol}_{i}\left(\pi^{\prime},\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}\right)=\phi_{i}\left(\pi^{\prime},\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}\right),$
$\Rightarrow\textsc{MargSol}_{i}\left(\pi^{\prime},v\right)=\phi_{i}\left(\pi^{\prime},v\right).$
The first implication holds by summing over all $\pi\in\Pi$. The second implication is by SA for both solution concepts. The third implication is by Equation˜5. Since $v$ was arbitrary, we conclude that $\phi$ is the same as MargSol. ∎
B.4 Proof of Theorem˜3
To prove this result, we first need the following lemma.
Lemma 6.
Consider a TCG $\langle N,u_{\pi,\alpha}\rangle$ for a given $\pi\in\Pi$ and $\alpha$, where $u_{\pi,\alpha}$ is defined as follows.
$u_{\pi,\alpha}(\pi^{\prime})=\begin{cases}\alpha,&\text{if }\pi\sqsubset\pi^{\prime}\\ 0,&\text{otherwise}\end{cases}$
If an extended solution concept $\psi\in\Psi$ satisfies EE and ENP, it must be true that
$\psi_{i}(u_{\pi,\alpha})=\begin{cases}\frac{(n-|P(\pi)|)!}{n!}\alpha,&\text{if }i=\ell(\pi)\\ 0,&\text{otherwise}\end{cases}$
Proof.
We know from Lemma˜4 that every $i\in N\setminus\{\ell(\pi)\}$ is a null player and should get zero reward share by any extended solution concept $\psi$ by ENP. Hence, the only positive reward should go to $\ell(\pi)$. By EE, we get
$\sum_{i\in P(\pi^{\prime})}\psi_{i}(u_{\pi,\alpha})=\frac{1}{n!}\sum_{\pi^{\prime}\in\Pi:P(\pi^{\prime})=N}u_{\pi,\alpha}(\pi^{\prime}).$
Given the earlier arguments, this reward will go entirely to $\ell(\pi)$, i.e., For $i$, the last player of $p$, we get that $\psi_{\ell(\pi)}(u_{\pi,\alpha})=\frac{1}{n!}\sum_{\pi^{\prime}\in\Pi:P(\pi^{\prime})=N}u_{\pi,\alpha}(\pi^{\prime})$.
By definition of $u_{\pi,\alpha}$, as long as $\pi$ is a prefix of a given sequence of players $\pi^{\prime}$, the remaining $N\setminus P(\pi)$ players can appear in any sequence. So, there are $(n-|P(\pi)|)!$ different $\pi^{\prime}\in\Pi$ such that $\pi$ is a prefix of $\pi^{\prime}$. Therefore, $u_{\pi,\alpha}(\pi^{\prime})=\alpha$ for all those $\pi^{\prime}$. Hence, $\psi_{\ell(\pi)}(u_{\pi,\alpha})=\frac{(n-|P(\pi)|)!}{n!}\alpha$. ∎
Now we present the proof of Theorem˜3.
Proof.
($\Leftarrow$) We have already proved in Theorem˜2 that MargSol satisfies SE, SA, and SNP, and by Lemma˜2, its reduced solution concept Ext-Shap satisfies EE, EA, and ENP.
($\Rightarrow$) This direction of the proof follows in similar lines to that of Theorem˜2. Assume that there exists an extended solution concept $\psi\neq\textsc{Ext-Shap}{}$ that also satisfies the three given properties.
For each carrier game, we know the structure of the unique solution concept that satisfies SE and SNP from Lemma˜6. Note that Ext-Shap satisfies all three properties for every game $\langle N,v\rangle$, hence it must satisfy them for the carrier games as well. Hence $\psi$ must be same as Ext-Shap for the carrier game $\langle N,u_{\pi,\alpha}\rangle$. Hence we have
$\textsc{Ext-Shap}_{i}(u_{\pi,\alpha_{\pi}})$ $=\psi_{i}(u_{\pi,\alpha_{\pi}}),\ \forall i\in N,\forall\pi\in\Pi,$
$\Rightarrow$ $\sum_{\pi\in\Pi}\textsc{Ext-Shap}_{i}(u_{\pi,\alpha_{\pi}})$ $=\sum_{\pi\in\Pi}\psi_{i}(u_{\pi,\alpha_{\pi}}),\ \forall i\in N,$
$\Rightarrow$ $\textsc{Ext-Shap}_{i}\left(\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}\right)$ $=\psi_{i}\left(\sum_{\pi\in\Pi}u_{\pi,\alpha_{\pi}}\right),\ \forall i\in N,$
$\Rightarrow$ $\textsc{Ext-Shap}_{i}\left(v\right)$ $=\psi_{i}\left(v\right),\ \forall i\in N.$
The first implication holds by summing over all $\pi\in\Pi$. The second implication is by EA for both solution concepts. The third implication is by Equation˜5. Since $v$ was arbitrary, we conclude that $\psi$ is the same as Ext-Shap. ∎
Appendix C Appendix to Section˜7
C.1 Proof of Lemma˜5
Proof.
Consider a solution concept $\phi\in\textsc{SeqShare}{}$ such that $\overline{\phi}=\textsc{Ext-Shap}$. Suppose $\exists i\in N$ such that $\phi_{i}(\pi^{*}(v))<\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi_{i}+i)-v(\pi_{i})$. Since $\phi\in\textsc{SeqShare}{}$, it satisfies I4OA, i.e.,
$\phi_{i}(\pi)\leqslant\phi_{i}(\pi^{*}(v)),\ \forall\pi\in\Pi$
$\implies$ $\phi_{i}(\pi)<\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi_{i}+i)-v(\pi_{i}),\ \forall\pi\in\Pi.$
Averaging over all possible $\pi\in\Pi$ such that $P(\pi)=N$, we get
$\overline{\phi}_{i}=$ $\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi)$ (6)
$<$ $\frac{\textstyle 1}{\textstyle n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi_{i}+i)-v(\pi_{i})=\textsc{Ext-Shap}_{i}.$ (7)
This contradicts the hypothesis that $\overline{\phi}=\textsc{Ext-Shap}$. ∎